RC meeting and Run Plan, Thursday Oct. 19, 2023

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information

Time: Monday - Friday at 3:30pm

  • Location: Counting house conference room
  • ZoomGov Meeting Connection Information


Useful Links


Hall A call-in instructions:

  • Call RC first.
  • Techs on site 7am-4:30pm Mon-Thu, 7-11am Fri. RC will call Hall A work coordinator (Zach) during normal working hours. After hours, RC calls Tech on-call.
  • After 9pm everything non-emergency needs to be deferred. What is emergency and what's not? Call RC. RC talks with Mark to decide.

Hall Access procedure:

  • Access check list (binder on the counting room desk) needs to be filled out before making an access.
  • Scan (take pictures with phone) and post it to the logbook.
  • Any work inside the SBS, BB boundary need to have PS locked out -- requires two QEWs to throw the big switch. Exception: field measurements.

New instruction for shift crew when changing target:

  • Beam restoration instructions have been updated. Printed version is in the counting room.
  • When we change target, make sure the SOFs of Target BPMs are set to the correct values.
  • SOFs for different targets: Needs an update after restoring beam tonight (Note: SOFs for major target positions are on the Control Room whiteboard.)

Status Update

Overall Summary

Owl Shift

  • Waiting for Beam most of the shift
  • Started and finished Moller measurement, 80.6 %

Day Shift Plan

  • Take single carbon foil sieve data with magnets OFF
  • Take multi carbon foil sieve data with magnets ON
  • Remove sieve slit with escorted access

Swing/Owl Shift Plan

  • Production Data


  • Beam down ~5 hours due to master oscillator issue during owl
  • Beam down 2.5+ hours for South Linac access

Opportunistically between spin ups (?):

  • H2 Reference data

Accelerator status

  • TBD
  • If you want to get into the hall, please communicate with the RC to get this organized!

Data collected :

0 C production up to 12 pm

Hall A Subsystems Status


  • Laser power and alignment were adjusted yesterday, spinning up well now


  • GEMs HV off in preparation for beam establishment in the hall.
  • BBCal 7.4 dead. Discussion on changing PMT base at next available time. (? still true?)


  • SBS and BB are on.

Hall access requests

If anyone wants to make an access, let the RC know!

Shift schedule


  • Friday 10/20 Day (Shift Lead)
  • Saturday 10/21 DAY (Target Operator, trade requested)

Please check when your shifts are!


Please sign up for any open shifts!

Running Plan

  • Standard production running at 45uA, with NMRs every 3 hours.

==> List of items in case He3 production running isn’t ready:

  1. Take data on reference cell at 10 atm H2 (portions that add up to 1 day) [85% (4200 elastic events) more data required for Bigbite momentum calibration]
  2. BigBite/single foil location reconstruction (~20 mins?) - BigBite + HCAL detector experts on call consulting analysis experts like Andrew/Sean (magnet on) (? is this different than the optics runs?)